OpenStudy (summersnow8):

Wheel of savings sale, regions are divided into 12 regions, four of those regions are red & award 5%. The three white award 10%, two blue regions award 25% discount, two green award 30% discount. The remaining gold region, gets a 50% discount. 1) what is the probability that exactly two of the first four customers get 10% discount?

OpenStudy (anonymous):

is ans = 0.21?

OpenStudy (lukebluefive):

First, the probability of any customer getting a 10% is the same as getting a white award. Since there are twelve regions, only three of which are white, the probability of landing in a white region is 3/12, or one out of four.

OpenStudy (lukebluefive):

Second, you multiply this probability by the number of times it needs to happen, namely twice. Therefore, this equals: \[\frac{1}{4} * \frac{1}{4} = \frac{1}{16}\]

OpenStudy (kropot72):

The probability of getting a 10% discount is 3/12. The binomial distribution will give the required probability as follows: \[P(2\ out\ of\ 4)=4C2 \times(0.25)^{2} \times (0.75)^{2}=6\times (0.25)^{2} \times (0.75)^{2}=you\ can\ calculate\]